Write an equation of the line that passes through the given point and is perpendicular to the given line.
problem 1.
(1,-1),y=3x+2
problem 2.
(5,0),y=2/3x-4
problem 3.
(3,-7),y=-1/5x+1

Answer by mananth(16946)

1 y = 3 x + 2
Divide by 1
y = 3.000 x + 2
Compare this equation with y=mx+b
slope m = 3

The slope of a line perpendicular to the above line will be the negative reciprocal m1*m2=-1
The slope of the required line will be -1/3

m= - 1/3 ,point ( 1 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = -1/3 + b
b= -2/3

m= - 1/3
The required equation is y = - 1/ 3 x - 2/ 3

----------------
1 y = 2/3 x -4
Divide by 1
y = 0.667 x + -4
Compare this equation with y=mx+b
slope m = 2/3

The slope of a line perpendicular to the above line will be the negative reciprocal -1.5
m1*m2=-1


m= -3/2 ,point ( 5 , 0 )
Find b by plugging the values of m & the point in
y=mx+b
0 = -15/2 + b
b= 15/2
m= -3/2
The required equation is y = -3/ 2 x + 15/ 2

m.ananth@hotmail.ca